OpenStudy (anonymous):

9. Determine if the lines are parallel, perpendicular or neither: y = -3/4x + 5 and 3x + 4y = -24 A. parallel B. perpendicular C. neither

OpenStudy (mathstudent55):

Parallel lines have the same slope. The slope-intercept form of the equation of a line is y = mx + b, where m = slope, and b = y-intercept. If you write both equations in this form, it will be easy to see if their slopes are the same. The first equation, y = -(3/4)x + 5, is already in y = mx + b form. Its slope is -3/4. Solve the second equatiuon for y, and see what m is equal to. 3x + 4y = 24 Subtract 3x from both sides: 4y = -3x + 24 Divide both sides by 4: y = -(3/4)x + 6 The slope of the second line is also -3/4. In addition, the y-intercepts are different, so you have two parallel lines.

OpenStudy (anonymous):

thanks

OpenStudy (mathstudent55):

wlcm